OpenStudy (anonymous):

Guiiiiseeee, so I have this question, please help? :,) The formula K equals C plus two hundred seventy-three shows how to estimate the temperature K measured in degrees Kelvin (°K) when you know the temperature C measured in degrees Celsius (°C). If room temperature is 295°K, what is room temperature in °C?

OpenStudy (anonymous):

K = C + 273 so C = K - 273 you are given K and you want C

OpenStudy (anonymous):

K=C+273

OpenStudy (anonymous):

None of those are the answe :(

OpenStudy (anonymous):

my options are, -22°C 22°C 568°C -568°C

OpenStudy (anonymous):

295-273

OpenStudy (anonymous):

c=k-273 => c=295-273

OpenStudy (anonymous):

=22